A basket contains red ball and 9 white balls. If four balls are taken from the basket one after the other, determine the possible values of the random variable W representing the number of white balls and construct the probability distribution of the random variable W

1
Expert's answer
2022-03-10T06:40:21-0500

Since there are 1 red ball, 9 white balls and four balls are selected, then W = {3, 4}

"P(W=3)={\frac {{9 \choose 3}*{1 \choose 1}} {{10 \choose 4}}}={\frac {84} {210}}=0.4"

"P(W=4)={\frac {{9 \choose 4}} {{10 \choose 4}}}={\frac {126} {210}}=0.6"